1.    If an individual’s preferences are described by the utility function U(X1 , X2 ) = X12 + X22,

a)    graph the indifference curve for U = 20 and U= 40.

b)    Find the optimal consumption quantities if P1 = US$2.50 ;                     P2 = US$ 7.50; and M = US$ 60.


1
Expert's answer
2021-06-17T12:39:33-0400

Part a

U(X1 , X2 ) = X12 + X22

The indifference curve is as shown below


If "P_1=2.5; P_2=2.5; m=60" , the budget line will be as follows

"P_1x_1+P_2x_2 =m"

"2.5x_1+2.5x_2 =60"

Also from U(X1 , X2 ) = X12 + X22

"MR_{(x_1,x_2)}=\frac{mUx_1}{mUx_2}" as "mUx_1 = 2x_1 \& mUx_2 = 2x_2"

So, "MR_{(x_1,x_2)}=\frac{2x_1}{2x_2} =\frac{x_1}{x_2}"


Part b

Equating the slope of budget line and MRs

"\frac{P_1}{P_2} =MRs_{(x_1,x_2)} \implies \frac{2.5}{7.5}=\frac{x_1}{x_2}"

"x_2=3x_1"

Therefore, "2.5x_1+7.5(3x_1)=60 \implies x_1= \frac{60}{25}=2.4; x_2= 7.2"